OpenStudy (anonymous):

square root of 72 x^2

OpenStudy (anonymous):

\[\sqrt{72x^2}\]

OpenStudy (anonymous):

Do you want a straight answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

anyways: \[\sqrt{2} * \sqrt{36x}\] if im not mistaken the answer is: \[6x\sqrt{2}\]

OpenStudy (anonymous):

Bump it up so someone else can check it just in case. Im pretty sure that's the answer though:)

OpenStudy (anonymous):

Thanks! What about 250h^4k^5?

OpenStudy (anonymous):

\[250h^4k^5\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I tried it but it confused me so try bumping it so someone else can help cuz I'm not sure:/